OpenStudy (anonymous):

which function defines the nth term of a the sequence 45,38,31,24,17,...

OpenStudy (solomonzelman):

the difference is ?

OpenStudy (anonymous):

A(n)=52+7n,A(n)=45-7n,A(n)=52-7n,or A(n)=46-n

OpenStudy (solomonzelman):

can you find the difference ?